### v1

# CrochetBench: Can Vision-Language Models Move from Describing to Doing in Crochet Domain?

Abstract
We present CrochetBench, a benchmark for evaluating the ability of multimodal large language models to perform fine-grained, low-level procedural reasoning in the domain of crochet. Unlike prior benchmarks that focus on high-level description or visual question answering, CrochetBench shifts the emphasis from describing to doing: models are required to recognize stitches, select structurally appropriate instructions, and generate compilable crochet procedures. We adopt the CrochetPARADE DSL as our intermediate representation, enabling structural validation and functional evaluation via execution. The benchmark covers tasks including stitch classification, instruction grounding, and both natural language and image-to-DSL translation. Across all tasks, performance sharply declines as the evaluation shifts from surface-level similarity to executable correctness, exposing limitations in long-range symbolic reasoning and 3D-aware procedural synthesis. CrochetBench offers a new lens for assessing procedural competence in multimodal models and highlights the gap between surface-level understanding and executable precision in real-world creative domains. Code is available at https://github.com/Peiyu-Georgia-Li/crochetBench.
1 Introduction
Procedural crafts such as crochet present a distinctive frontier for multimodal learning. Unlike traditional captioning or recipe datasets (chefFusion; Hu_2023_ICCV; mohbat2024llava), crochet patterns intertwine three interdependent modalities: (i) structured symbolic language, where stitch abbreviations and counts define a precise grammar of construction; (ii) long-form natural language, which provides contextual guidance such as materials and sizing; and (iii) visual evidence, including photographs of completed objects and motif diagrams. Success requires not just alignment across modalities but step-wise reasoning that preserves procedural fidelity, making the challenge closer to program synthesis than generic description.
Crochet also offers a unique testbed for 3D-aware reasoning. Each stitch encodes both local geometry and global connectivity, forming a topological structure that must be preserved across steps. Generating or interpreting patterns thus demands reasoning over how sequential operations accumulate into volumetric form. In effect, crochet couples symbolic instruction following with embodied spatial reasoning, cultivating abilities essential for domains where language must ground into physical tasks.
Despite the rapid growth of multimodal benchmarks (fu2024mmecomprehensiveevaluationbenchmark; li2023seed; zhang-etal-2025-ngqa; yue2024mmmu), existing datasets have largely focused on description or grounding. COCO (lin2015microsoftcococommonobjects) catalyzed captioning research, TextCaps (sidorov2020textcapsdatasetimagecaptioning) extended it to text-in-the-wild, and Recipe1M (marin2019recipe1mdatasetlearningcrossmodal) explored cross-modal cooking instructions. While recipes also involve multi-step procedures, validating correctness typically requires real-world execution, making large-scale evaluation slow and resource-intensive. Crochet, by contrast, provides a symbolic domain where outputs can be automatically verified through DSL compilation, enabling scalable and efficient study of step-wise reasoning. For a more detailed survey, see Appendix A. Yet these benchmarks stop short of testing whether models can follow symbolic grammars, respect numerical and spatial constraints, and produce outputs that are executable. Current systems can describe, but not reliably do.
CrochetBench fills this gap by centering evaluation on instructional fidelity: can models not only recognize and generate, but also output step-wise, compilable instructions that respect symbolic, numerical, and topological structure? Each example in CrochetBench is a multimodal package—structured JSON metadata (stitch inventories and abbreviations), full-text procedures with rows/rounds and conditionals, and paired images of finished objects and motifs. Crucially, CrochetBench is paired with CrochetPARADE (crochetparade), a domain-specific language (DSL) enabling executable evaluation, where natural language instructions are translated into compilable code enforcing geometric and topological coherence.
Our contributions are fourfold: (1) CrochetBench, the first executable benchmark for procedural crafts, unifying symbolic, textual, and visual modalities with evaluation protocols emphasizing procedural fidelity and 3D-aware reasoning; (2) a comprehensive task suite spanning recognition, comprehension, generation, and DSL translation; (3) integration of CrochetPARADE into an executable pipeline, enabling scalable, automated verification of outputs—unlike domains such as cooking, which require real-world execution—thereby shifting evaluation from surface similarity to procedural fidelity; and (4) baseline analyses of state-of-the-art VLMs/MLLMs, revealing systematic weaknesses including hallucinations, captioning bias, and structural artifacts.
Taken together, CrochetBench opens a new direction for multimodal research: moving beyond describing what we see, toward generating executable procedures that respect symbolic grammar, numerical accuracy, and topological coherence—paving the way for models that can reason in structured 3D spaces.
2 Dataset Description
Figure 1: Distribution of the top-10 most common project types in CrochetBench.
CrochetBench is a large-scale, structured benchmark consisting of 6,085 crochet patterns spanning 55 distinct project categories. It was constructed by collecting publicly available patterns from the Yarnspirations website [^1] https://www.yarnspirations.com/collections/patterns, a widely used repository for fiber arts. The raw patterns were originally formatted as PDF documents, which were parsed and normalized using a GPT-4o-mini-based conversion pipeline. This process extracted and standardized key fields including pattern metadata, materials, measurements, gauge, abbreviations, and full step-by-step instructions. Each entry was converted into a machine-readable JSON object with a consistent schema, and 98.77% of patterns include an associated product image, enabling multimodal supervision.
Figure 2: Skill level distribution across the CrochetBench dataset. Note that the “Experienced” slice (1.3%) is annotated externally due to its small size.
The dataset supports diverse real-world crochet practices, with project types ranging from simple accessories to complex garments. Figure 1 lists the ten most common categories by frequency. The majority of patterns belong to a small number of dominant types—Afghans and Blankets alone account for over one-quarter of the dataset. More details can be found at Appendix B.
Each pattern is labeled with one of four primary skill levels—beginner, easy, intermediate, or experienced. This allows for stratified evaluation across complexity tiers. Figure 2 shows the skill level distribution, which is strongly skewed toward beginner-friendly content. Only one pattern (0.02%) is missing a skill level label. More details can be found at Appendix B.
Instructional complexity varies substantially across patterns. The number of characters in each instruction ranges from 20 to over 30,000, with a mean of 3,216 and a median of 2,453. Abbreviation counts (i.e., unique stitch tokens per pattern) range from 1 to 31, with an average of 10.6. These statistics are summarized in Appendix B.1. We observe a clear correlation between skill level and instruction length: beginner patterns tend to be short and use fewer abbreviations, while experienced patterns are significantly longer and more symbolically dense.
In addition to symbolic complexity, the dataset contains 3,143 abbreviation instances mapped to 789 unique standardized stitch tokens. This lexical mapping enables tasks such as vocabulary translation, sequence generation, and instruction validation. Beyond raw instructions, the structured schema also records rich metadata, including gauge, hook size, yarn weight, and measurements. A representative dataset entry is provided in Appendix 10.
Overall, CrochetBench provides a rich resource for multimodal modeling, symbolic reasoning, and structure-aware generation. Its coverage across diverse categories and complexity levels enables broad benchmarking of both open-ended generation and instruction fidelity tasks.
Table 1: Overall statistics of the CrochetBench dataset.
Total Patterns Image Coverage Avg. Instr. Length #Project Types
CrochetBench 6,085 98.77% 3,216 characters 55
3 Experiment
We empirically evaluate CrochetBench by defining a set of structured tasks and benchmarking a diverse pool of multimodal large language models. Our experiments are designed to test models across progressively challenging stages, from low-level recognition to high-level executable synthesis.
Benchmark Tasks.
Table 2 summarizes the four evaluation tasks in CrochetBench, which progress systematically from recognition to comprehension, generation, and ultimately executable synthesis. Task A (Stitch Recognition) evaluates a model’s ability to detect symbolic primitives in crochet images, establishing the foundation for multimodal perception. Task B (Instruction Selection) requires models to align visual evidence with candidate textual instructions, thereby testing multimodal grounding and fine-grained comprehension. Unlike conventional description tasks, the candidates are procedural steps rather than captions, and correct selection often requires reasoning about how local steps contribute to the final product. Task C (Instruction Generation) advances from comprehension to open-ended production, challenging models to generate natural-language procedural instructions that are both perceptually grounded and linguistically faithful to domain conventions. Evaluation here emphasizes lexical and symbolic fidelity, but does not directly capture structural validity. Task D (Instruction-to-DSL Translation) addresses this gap by requiring models to output a compilable program in a domain-specific language (DSL). The step-level variant tests local semantic grounding, while the project-level variant demands global structural consistency across the entire pattern. Compilation-based evaluation directly measures executable faithfulness, ensuring that generated instructions are not only linguistically plausible but also structurally sound. Together, these tasks form a structured evaluation ladder that moves from perceptual recognition to programmatic execution, probing both low-level perception and high-level symbolic reasoning.
Table 2: Overview of benchmark tasks in CrochetBench. Tasks progress from recognition to comprehension, generation, and executable synthesis.
ID Ability Tested Task Evaluation Metrics Test Size
A Recognition Stitch Recognition F1, Precision, Recall
6,009
(CrochetBench-A)
B Comprehension Instruction Selection Accuracy
6,003
(CrochetBench-B)
C Generation Instruction Generation BLEU, ROUGE, ChrF
6,009
(CrochetBench-C)
D Formalization Instr.-to-DSL (Step) Valid Pattern Rate
119
(CrochetBench-Dstep)
Instr.-to-DSL (Project) Valid Pattern Rate
100
(CrochetBench-Dproj)
Model Selection.
We evaluate a diverse set of vision–language models (VLMs) spanning both open-source and closed-source families. The open-source group includes Salesforce BLIP-2 Flan-T5 XL (3B), a perception-focused baseline widely used in image–text tasks; Google Gemma 3 (4B) and Qwen2-VL (7B), two recent models trained with large-scale multimodal alignment; and DeepSeek-VL (7B), a larger open-source model designed for stronger vision–language reasoning. On the closed-source side, we assess GPT-4o, Gemini 2.5 Flash-Lite, and Claude Sonnet 4, which represent state-of-the-art commercial VLMs that push the limits of multimodal reasoning. This selection spans perception-heavy baselines to general-purpose multimodal reasoning systems, enabling us to evaluate capabilities across both research-grade and production-grade settings. [^2] We use the term “VLM” broadly to include both traditional vision–language models (e.g., BLIP-2) and modern multimodal large language models (e.g., GPT-4o, Gemini, Claude).
3.1 Task A: Stitch Recognition
The first task evaluates a model’s ability to identify crochet stitch types from an image of a finished product. We construct CrochetBench-A, a subset of 6,009 examples from the full benchmark, where each product image is paired with ground-truth stitch annotations. These labels are derived from the official pattern instructions and normalized into a standardized set of stitch abbreviations (e.g., sc, hdc, dc) to ensure consistency across patterns. Unlike standard image classification, this is a multi-label prediction problem: multiple stitches may co-occur within the same image, often with subtle visual differences in texture and geometry. This task therefore probes fine-grained visual grounding of structured crochet semantics.
Evaluation.
For each example, we compute overlap between the predicted and reference stitch sets. True Positives (TP) are stitches correctly predicted; False Positives (FP) are stitches predicted but not in the reference; and False Negatives (FN) are stitches in the reference but missed by the model. From these counts, we compute precision (fraction of correct predictions among all predictions), recall (fraction of ground-truth stitches recovered), and F1 score (harmonic mean). Metrics are averaged across examples to provide overall performance. This formulation rewards models that recover all present stitches while avoiding spurious predictions.
Accurate stitch recognition is foundational for the benchmark, as later tasks (e.g., instruction selection and instruction generation) depend on robust detection of stitch primitives.
Table 3: Evaluation results on the Stitch Recognition task. We report Precision, Recall, and F1. Best results are bold; second-best are underlined.
Model Size Precision Recall F1
Open Source Salesforce BLIP-2 Flan-T5 XL 3B 0.2953 0.2303 0.2250
Google Gemma 3 4B 0.2054 0.1021 0.1265
DeepSeek-VL 7B 0.5447 0.7476 0.6060
Qwen2-VL 7B 0.5414 0.6974 0.5816
Closed Source GPT-4o 0.6214 0.5939 0.5801
Gemini 2.5 Flash-Lite 0.7449 0.4977 0.5683
Claude Sonnet 4 0.7861 0.5312 0.6094
Results. Claude Sonnet 4 achieves the best overall F1 score, demonstrating strong precision in stitch recognition. Among open-source models, DeepSeek-VL performs best, with notably high recall, while Qwen2-VL offers a competitive balance. These results highlight the gap between commercial VLMs and open-source alternatives, but also suggest that large open-source models are beginning to approach closed-source performance in fine-grained recognition.
3.2 Task B: Instruction Selection
The second task evaluates whether a model can correctly align a finished crochet product image with its corresponding natural-language instruction. We construct CrochetBench-B, a subset of 6,003 examples, where each item consists of one correct instruction and three distractor instructions sampled from the same project category (e.g., hats, rugs). Distractors are carefully chosen to share structural motifs and vocabulary, thereby increasing difficulty and requiring fine-grained visual–textual alignment rather than reliance on superficial cues. The answer distribution across options is approximately uniform (A: 24.9%, B: 25.7%, C: 23.7%, D: 25.7%).
Evaluation.
To enable scalable assessment, we formulate the task as a 4-way multiple-choice question (MCQ). The model must select one option (A–D), with exactly one correct answer. Predictions are extracted using a lightweight regex-based method that identifies explicit letter outputs (e.g., “A”, “The answer is B”). If no parsable choice is found, the response is marked as unanswered. Accuracy is reported as the evaluation metric.
This task provides a controlled measure of visual grounding and semantic alignment between image content and procedural text, without requiring generative modeling. It highlights the challenge of distinguishing between subtle visual cues and domain-specific terminology, which is essential for bridging perception and structured instruction understanding.
Table 4: Evaluation results on the Instruction Selection task (4-way multiple choice). We report Accuracy. Best results are bold; second-best are underlined.
Model Size Accuracy
Open Source Salesforce BLIP-2 Flan-T5 XL 3B 0.2562
Google Gemma 3 4B 0.2494
DeepSeek-VL 7B 0.2892
Qwen2-VL 7B 0.4196
Closed Source GPT-4o 0.5811
Gemini 2.5 Flash-Lite 0.5563
Claude Sonnet 4 0.5739
Results. GPT-4o achieves the highest accuracy, closely followed by Claude and Gemini, all of which substantially outperform open-source VLMs. Among open-source models, Qwen2-VL is the strongest, while BLIP-2 and Gemma remain near chance level, underscoring the difficulty of visually grounded instruction matching in this domain.
3.3 Task C: Instruction Generation
The third task evaluates a model’s ability to generate natural-language crochet instructions directly from an image of a finished product. We construct CrochetBench-C, a subset of 6,009 examples, where each product image is paired with its full pattern. Unlike classification or selection, this task requires producing a multi-step sequence that follows domain-specific syntax, ordering, and stitch logic. Outputs are expected to resemble real-world crochet patterns, written line by line (e.g., “Rnd 1: ch 4, 6 sc in ring”). This setting challenges models to translate visual evidence into coherent procedural text that preserves both symbolic accuracy and structural consistency.
Evaluation.
We evaluate generation quality using standard text-generation metrics and a domain-specific structural measure. BLEU and ROUGE-L capture lexical and n-gram overlap with the reference instructions. ChrF, computed over character n-grams, provides a finer-grained signal of similarity and is particularly suited to stitch abbreviations. While these metrics measure surface-level overlap, they do not directly test whether the generated instructions are structurally consistent (e.g., balanced stitch counts, valid round progression). To address this limitation, we complement Task C with Task D, where compilation-based evaluation of DSL translations directly assesses executable consistency. Together, these tasks allow us to probe both linguistic fidelity and procedural correctness.
Table 5: Evaluation results on the Instruction Generation task. We report BLEU, ROUGE-L, and ChrF. Higher is better. Best results are bold; second-best are underlined.
Model Size BLEU ROUGE-L ChrF
Open Source Salesforce BLIP-2 Flan-T5 XL 3B 0.0021 0.0926 9.32
Google Gemma 3 4B 0.0010 0.0329 5.17
Qwen2-VL 7B 0.0160 0.2084 15.76
DeepSeek-VL 7B 0.0133 0.1968 18.12
Closed Source GPT-4o 0.0333 0.2353 23.80
Gemini 2.5 Flash-Lite 0.0482 0.2583 30.20
Claude Sonnet 4 0.0331 0.2516 22.95
Results. Gemini achieves the strongest overall performance across all metrics, substantially outperforming other closed-source models. GPT-4o and Claude follow closely, while open-source models lag behind with significantly lower BLEU and ChrF scores. This gap highlights the difficulty of generating structurally faithful crochet instructions, which requires models to capture both visual details and domain-specific procedural logic.
3.4 Task D: Instruction-to-DSL Translation
We construct two subsets for Task D: CrochetBench-Dstep (119 items) and CrochetBench-Dproj (100 items). For CrochetBench-Dstep, we manually annotate the correct CrochetPARADE DSL for the previous context of each pattern (a prefix of NL (natural language) –DSL pairs). The model is then given this prefix along with the next natural-language instruction, and must predict the corresponding DSL line. Predictions are evaluated by checking compilation validity with the CrochetPARADE validator. For CrochetBench-Dproj, we adopt a few-shot prompting setup rather than full manual annotation. Annotators provide a single reference program in CrochetPARADE as context, which is paired with a new natural-language instruction and its product image. The model must then generate a complete DSL program, and outputs are assessed by whether they compile fully or partially using the validator.
Design Rationale. Crochet patterns admit many valid DSL realizations (e.g., alternative groupings or equivalent constructs), meaning that no single gold reference is canonical. A reference-based metric could unfairly penalize models that produce semantically correct but structurally different programs. By instead relying on the validator, CrochetBench shifts evaluation away from surface-level string matching toward functional executability—the central criterion for crochet synthesis. This design is consistent with other program synthesis benchmarks, such as semantic parsing and SQL generation, where execution accuracy is often preferred over exact string match.
3.4.1 Step-Level Translation
In the step-level setting, the model receives a prefix of NL–DSL pairs and must generate the DSL line corresponding to the next natural-language instruction. This setup reflects an incremental synthesis process in which correctness depends on maintaining stitch-level consistency across steps. Since crochet patterns are inherently stateful, earlier context is critical for resolving constructs such as increases, repeats, and turning chains. To capture progression through a pattern, we sample 52 early examples (steps 1–2), 34 mid examples (steps 3–4), and 33 late examples (steps 5–6). Step-level inputs are formatted as:
$\text{Prefix (NL--DSL pairs)}+\text{Next NL instruction}\;\;\rightarrow\;\;\text{Next DSL line}.$
This formulation allows models to generate locally plausible DSL steps, but global correctness ultimately requires stronger contextual reasoning.
Evaluation.
We evaluate models using Compilation Success Rate (CSR), defined as the proportion of generated DSL outputs that compile successfully with the CrochetPARADE validator. This validator-based metric is a key strength of CrochetBench: unlike reference-based string matching, CSR directly measures whether generated programs are executable, rewarding functional correctness rather than surface similarity. To better understand systematic errors, we also conduct fine-grained error analysis across four categories: (1) syntax structure errors, (2) stitch definition errors, (3) labeling and reference errors, and (4) structural or formatting issues. A detailed taxonomy with illustrative examples is provided in the Appendix.
CrochetBench further supports multimodal verification, as CrochetPARADE programs can be rendered into simulated product images. These renderings can be compared with ground-truth product images using pretrained vision–language models (e.g., CLIP), enabling evaluation of both structural validity and visual faithfulness. While current model performance leaves ample room for progress, the combination of validator-based execution checks and prospective image-grounded verification establishes CrochetBench as a uniquely rigorous testbed for structured multimodal reasoning.
Table 6: Verification results across models, grouped by category.
Model CSR (%) Undef. Br. MRef Other Tot. Err.
Open Source
Salesforce BLIP-2 Flan-T5 XL 4.2 29.8 48.2 20.2 1.8 114
Google Gemma 3 3.4 26.1 10.4 63.5 115
DeepSeek-VL 32.8 36.2 38.8 11.2 13.8 80
Qwen2-VL 35.3 42.9 39.0 1.3 16.8 77
Closed Source
GPT-4o 36.1 39.5 43.4 17.1 76
Gemini 2.5 Flash-Lite 41.2 51.4 21.4 5.7 21.5 70
Claude Sonnet 4 52.1 45.6 28.1 26.3 57
Results. Step-level translation is locally tractable: models often produce syntactically valid lines, but compilation failures reveal semantic underspecification. Claude achieves the highest valid rate (52.1%), while DeepSeek-VL and Qwen2-VL are the strongest among open-source models. Errors are dominated by syntax (brackets, references) and undefined stitches, underscoring the difficulty of maintaining state consistency even when local syntax is correct.
3.4.2 Project-Level Translation
In the project-level setting, the model is provided with the complete crochet instruction in natural language together with the corresponding product image, and must generate an entire CrochetPARADE program. This variant is globally self-contained but considerably more challenging than the step-level task: models must track stitch states over long horizons, resolve ambiguities in natural language, and produce code that is both syntactically valid and semantically aligned with the final design. Image grounding plays a crucial role in disambiguating constructs such as repeated motifs, symmetry, and termination conditions.
Evaluation.
We assess model outputs using two complementary metrics. The first is Compilation Success Rate (CSR), and the second is Partial Executable Rate (PER), which measures the average fraction of a program that compiles successfully before failure. While CSR captures all-or-nothing executability, PER provides a finer-grained view of structural alignment, offering credit to models that generate correct prefixes even if the full program does not compile. Error types are categorized using the same taxonomy as in the step-level evaluation.
Table 7: Project-level verification results. We report Compilation Success Rate (CSR) and Partial Executable Rate (PER). Best scores are bold, second-best are underlined.
Model Size CSR (%) PER (%)
Open Source Salesforce BLIP-2 Flan-T5 XL 3B 1.0 0.00
Google Gemma 3 4B 1.6 5.29
DeepSeek-VL 7B 8.1 37.49
Qwen2-VL 7B 21.0 30.28
Closed Source GPT-4o 4.0 2.76
Gemini 2.5 Flash-Lite 4.0 5.67
Claude Sonnet 4 5.0 8.16
Table 8: Project-level error type distribution across models (percentage). Abbreviations: Undef. (Undefined), Br. (Brackets), Lbl. (Labels Missing), Non-adj. (Non-adjacent), MRef (Multi-Reference), Other (Other/Runtime).
Model Undef. Br. Lbl. Non-adj. MRef Other
Open Source
Salesforce BLIP-2 Flan-T5 XL 2.0 37.4 58.6 2.0
Google Gemma 3 39.3 14.8 11.5 8.2 14.8 11.5
DeepSeek-VL 36.3 45.1 4.4 11.0 3.3
Qwen2-VL 25.0 12.5 7.8 51.6 3.1
Closed Source
GPT-4o 61.5 7.3 16.7 4.2 4.2 6.2
Gemini 2.5 Flash-Lite 51.1 18.1 9.6 8.5 4.3 8.6
Claude Sonnet 4 46.3 14.7 21.1 7.4 5.3 5.3
Results. Project-level translation remains highly challenging. Qwen2-VL achieves the best valid rate (21.0%) and strong PER, surpassing all closed-source systems. DeepSeek-VL demonstrates robust partial executability despite a lower valid rate. By contrast, GPT-4o, Gemini, and Claude achieve lower scores, highlighting that even state-of-the-art VLMs struggle with long-range structural consistency in executable synthesis. Error analysis (Table 8) reveals that closed-source models often fail due to label and reference inconsistencies, while open-source models more frequently exhibit undefined stitches and syntax errors.
4 Discussion
The results across CrochetBench highlight both the promise and current limitations of Visual language models in bridging perception, procedural reasoning, and executable synthesis. By structuring tasks in a progressive ladder, we expose clear gradients of difficulty: while contemporary models demonstrate competence in low-level recognition and mid-level comprehension, their performance declines substantially when asked to generate or formalize instructions into compilable domain-specific programs. This pattern underscores both methodological bottlenecks in multimodal grounding and fundamental challenges in symbolic reasoning over long-horizon structures.
In Task A, stitch recognition is feasible for both open- and closed-source models, though precision–recall tradeoffs differ. Claude Sonnet 4 prioritizes precision for a higher F1, while DeepSeek-VL favors recall via richer visual encoders. This reflects divergent inductive biases: commercial models better regularize spurious predictions, while open-source models overpredict to capture subtle textures. Crucially, this is a multi-label, fine-grained texture classification task with symbolic implications, distinguishing it from generic object recognition. Baselines like BLIP-2, designed for captioning, underperform due to insufficient symbolic grounding (blip; clip). Task B amplifies these challenges. Closed-source models, especially GPT-4o, outperform open models by a wide margin, underscoring the difficulty of aligning visual cues with procedural semantics. Distractor instructions are intentionally plausible, demanding reasoning over local–global coherence, not just token overlap (GVE; VQA; miech2019howto100mlearningtextvideoembedding). Qwen2-VL’s competitive performance suggests that scale and pretraining diversity help, but commercial systems benefit from stronger instruction tuning and better vision–language alignment.
Task C exposes the steep drop in open-ended procedural generation. Gemini 2.5 leads across BLEU, ROUGE, and ChrF (papineni2002bleu; lin2004rouge; popovic2015chrf), reflecting fluency in structured text. Yet absolute scores are low, and outputs often fail to conform to crochet-specific syntax. Open models hallucinate frequently, lacking exposure to domain-aligned distributions. This highlights a broader issue: fluency in general language modeling does not imply competence in structured, domain-specific generation (miech2019howto100mlearningtextvideoembedding; alayrac2022flamingovisuallanguagemodel).
Task D proves most demanding. Even top closed-source models rarely exceed 6% project-level compilation success, while Qwen2-VL unexpectedly achieves 21%, suggesting better symbolic generalization under execution constraints. Syntax errors (e.g., undefined stitches, malformed brackets) dominate in open models, while closed models produce syntactically valid but semantically inconsistent programs. These complementary failure modes hint at differing generalization paths: symbolic robustness versus lexical fluency. Notably, Task D also tests 3D-aware procedural fidelity—models must translate visual or textual cues into symbolic programs that unfold into spatially coherent structures, not just grammatically valid sequences.
Taken together, these results demonstrate that success in linguistic generation does not translate directly to executable synthesis. Performance decays sharply when evaluation moves from surface-level fidelity (BLEU, ROUGE) to structural validity (compilation), reinforcing the importance of execution-grounded metrics for procedural tasks. The relative strength of Qwen2-VL at the project level further suggests that progress may come not from scaling alone, but from architectural or training adjustments that better capture long-range dependencies and stateful operations. More broadly, CrochetBench exposes a critical frontier for multimodal reasoning: models must not only ground text in vision but also internalize procedural invariants that guarantee functional correctness.
However, our analysis must be contextualized within certain constraints. The benchmarks reflect standardized stitch sets and normalized instructions, which, while ensuring comparability, simplify the variability encountered in real-world crochet practice (e.g., designer-specific shorthand, unconventional repeats). Moreover, evaluation metrics in Task C rely on string overlap, which may underestimate semantically correct but lexically divergent outputs. Even the compilation-based metric in Task D, though stronger, cannot assess visual fidelity unless paired with image-render verification. Finally, the sample size for project-level DSL translation remains limited, and absolute validity rates are low, constraining fine-grained statistical comparisons.
5 Future Work
A central avenue for future research is advancing models that translate free-form natural language crochet instructions into the formal CrochetPARADE DSL. As a domain-specific programming language, CrochetPARADE not only enables executable verification of multimodal outputs but also positions crochet as a program synthesis problem, where compilers map symbolic grammars into machine-executable instructions, extending ideas from prior DSL work in domains such as knitting (knitscript), graphics (ellis2018learninginfergraphicsprograms), and robotics (mu2024robocodexmultimodalcodegeneration). This perspective connects naturally to CAD/CAM integration: industrial crochet and warp-knitting machines (e.g., COMEZ, Jakob Müller) already rely on pipelines from graphical design interfaces, to intermediate graph-based representations, to low-level machine code. CrochetPARADE could serve as a standardized intermediate layer in this workflow, bridging human-facing authoring tools with machine-facing execution systems (text2cad).
Several technical directions arise from this framing. First, incorporating explicit state-tracking mechanisms—whether through memory-augmented architectures (graves2016hybrid) or symbolic scaffolds (nye2021workscratchpadsintermediatecomputation)—could mitigate long-range inconsistencies in DSL translation. Second, multimodal pretraining enriched with procedural and topological domains (e.g., assembly instructions (cao2023instruction), instructional videos (miech2019howto100mlearningtextvideoembedding)) may narrow the gap between natural-language fluency and executable synthesis. Third, hybrid evaluation pipelines that combine compilation checks with visual render comparisons could more holistically assess structural and perceptual fidelity, building on metrics used in programmatic 3D generation (jain2022zero; paschalidou2021neuralpartslearningexpressive). Finally, CrochetBench offers a testbed for exploring neuro-symbolic integration, where neural perception is paired with symbolic reasoning to reconcile fine-grained visual cues with globally coherent program execution.
In sum, CrochetBench and CrochetPARADE highlight a pronounced frontier at the intersection of multimodal learning, program synthesis, and digital fabrication. Progress here may one day close the loop from human intent to automated textile production, advancing both structured multimodal reasoning and computational craft.
Ethics Statement
We acknowledge that the original crochet pattern PDFs are protected under copyright and therefore do not distribute raw files or full texts. Instead, we release only structured JSON annotations generated with GPT, reference URLs to the original sources, and our parsing and annotation scripts. The benchmark is provided strictly for non-commercial academic use. This approach enables reproducible research while respecting intellectual property and ensuring that our dataset serves as a tool for studying structured generation rather than redistributing creative works.
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our results. All datasets, task templates, and evaluation procedures are documented in the main text and appendix. An anonymous repository containing the full source code, experiment scripts, and detailed reproduction instructions has been made publicly available at: https://github.com/Peiyu-Georgia-Li/crochetBench. This ensures that all reported results can be independently verified and extended by the research community.
Appendix A Related Work
A.1 Multi-modal Datasets Beyond Captioning
Most multimodal benchmarks have centered on descriptive pairing of images and natural language. Large-scale resources such as COCO [lin2015microsoftcococommonobjects] and Flickr30k [plummer2016flickr30kentitiescollectingregiontophrase] provide dense captions of everyday scenes, advancing vision–language representation learning. More recent datasets extend beyond captioning to procedural or instructional domains. Recipe1M+ aligns food images with ingredient lists and cooking steps [marin2019recipe1mdatasetlearningcrossmodal], while instructional video corpora such as YouCook2 [zhou2017automaticlearningproceduresweb] and HowTo100M [miech2019howto100mlearningtextvideoembedding] pair narrated demonstrations with visual segments. These resources emphasize semantic alignment but generally evaluate with retrieval- or similarity-based metrics.
Our benchmark departs from this paradigm by pairing images with executable procedures. Rather than asking models to generate a semantically similar description, we require them to synthesize a program (CrochetPARADE DSL) that can be rendered and structurally verified. This shift enables functional evaluation—akin to program synthesis—and reduces reliance on subjective similarity measures.
A.2 Bridging Visual Reasoning and Procedural Language
Procedural understanding benchmarks highlight the importance of sequential, state-dependent reasoning. Datasets such as Assembly101 [sener2022assembly101largescalemultiviewvideo] and IKEA-ASM [benshabat2023ikeaasmdatasetunderstanding] capture human assembly activities, modeling dependencies across actions, objects, and preconditions. Instructional video benchmarks further test long-horizon understanding and error detection [wei2025instructionbenchinstructionalvideounderstanding, lin2024videoguibenchmarkguiautomation]. Our task complements this line of work by grounding supervision not in temporally segmented actions but in artifact-centric procedures—finished crochet items paired with symbolic, stepwise instructions. This enables models to reason about topology, geometry, and sequential dependencies in a single unified representation.
A.3 Crafting and Domain-Specific Procedural Data
Closer to our domain, prior work has begun to explore fiber crafts. Seitz et al. introduced Digital Crochet, a visual, graph-based notation system for representing crochet patterns [digital_crochet]. In knitting, Kaspar et al. developed Neural Inverse Knitting, mapping images of knitted patterns to machine instructions [kaspar2019neuralinverseknittingimages]. These works underscore the feasibility of executable supervision in crafts, but remain limited in scale and scope. Our benchmark builds on this foundation by providing thousands of real, community-tested crochet patterns, along with a compiler and renderer for executable evaluation.
A.4 Relation to Code Benchmarks
The executable nature of CrochetPARADE connects it directly to program synthesis benchmarks. HumanEval [chen2021evaluatinglargelanguagemodels], MBPP [austin2021programsynthesislargelanguage], and APPS [hendrycks2021measuringcodingchallengecompetence] evaluate code generation by execution against unit tests. Similarly, Spider [yu2019spiderlargescalehumanlabeleddataset] and CoNaLa [yin2018mining] frame natural language to code translation tasks. In vision, Im2LaTeX-100K [deng2017imagetomarkupgenerationcoarsetofineattention] and pix2code [beltramelli2017pix2codegeneratingcodegraphical] evaluate image-to-program translation with render fidelity as the metric. Our tasks—ranging from text-to-DSL translation to image-to-stitch recognition—extend this paradigm to the crafting domain, introducing structured 2D/3D topology as part of the evaluation. This grounds multimodal learning in a setting where success requires both semantic alignment and structural correctness.
A.5 Domain-Specific Languages for Executable Evaluation
A key enabler of our benchmark is the use of a domain-specific language. Prior DSL-based benchmarks, such as Im2LaTeX and the CROHME competition on handwritten math recognition [crohme], demonstrate how symbolic formalisms enable reproducible rendering and structural evaluation. CrochetPARADE adopts this principle for fiber crafts: each pattern compiles into an abstract syntax tree that can be rendered and tested, supporting metrics such as ExecPass@K, structural unit tests, and image–render similarity. This functional perspective moves beyond surface-level similarity to test whether a model’s output actually works, in contrast to chemistry and recipe domain, which require quite amount of time to validate it [huang2024application, chefFusion, huang2025chemhgnn].
Appendix B Additional Dataset Statistics
B.1 Instruction Complexity by Skill Level
Skill Level Avg. Length Median Length Avg. Abbr. Count
Beginner 1,674 1,365 9.2 465
Easy 2,761 2,182 10.8 3,569
Intermediate 4,221 3,387 10.7 1,967
Experienced 7,689 6,729 9.8 80
Table 9: Instruction complexity by skill level. Length is measured in characters.
Example Dataset Entry
Table 10: Representative pattern entry from CrochetBench.
Field Value
Pattern Name SKULL TRICK OR TREAT BAG (TO CROCHET)
Skill Level Intermediate
Project Type Bags or Purses
Measurements 15 cm diameter $\times$ 15 cm high (excluding handle)
Gauge 13 sc and 14 rows = 10 cm
Materials Lily® Sugar’n Cream (White, Black), 5 mm hook, cardboard
Image https://www.yarnspirations.com/cdn/shop/products/SCC0303-005314M.jpg
Source input_file/Bags+Purses/SCC0303-005314M.pdf
Instructions (truncated for brevity)
B.2 Skill Level Distribution
B.2.1 Overall Distribution
Table 11 summarizes the overall distribution of skill levels across the CrochetBench dataset. The majority of patterns are labeled as easy (58.7%), followed by intermediate (32.3%). Only a small fraction are classified as beginner (7.6%) or experienced (1.3%). [^3] Three additional rare labels were observed: easy to intermediate (1 pattern), beginners (1 pattern), and beginner/easy (1 pattern). Together they account for $0.1\%$ of the dataset.
Table 11: Overall skill level distribution. Percentages are relative to all patterns with annotated skill levels.
Skill Level Count Percentage
Easy 3569 58.66%
Intermediate 1967 32.33%
Beginner 465 7.64%
Experienced 80 1.31%
Total 6084 100%
One pattern (0.02%) is missing an annotated skill level.
B.2.2 Distribution by Project Type
We further break down skill levels by the top 10 most common project types. Results are shown in Table 12. In most categories, easy patterns dominate, typically ranging between 53–70%. Intermediate is the second most common, while beginner and experienced remain consistently low across categories.
Table 12: Skill level distribution by top 10 project types. Percentages are within each project category.
Project Type Easy Intermediate Beginner Experienced
Afghans & Blankets 56.1% 35.3% 7.0% 1.5%
Hats 61.3% 27.8% 10.1% 0.7%
Sweaters & Cardigans 56.6% 35.9% 5.0% 2.5%
Shawls & Wraps 52.7% 41.8% 4.2% 1.2%
Scarves 63.2% 20.7% 16.1%
Pillows & Poufs 70.0% 22.9% 6.5% 0.7%
Amigurumi & Toys 64.0% 33.2% 2.1% 0.7%
Bags & Purses 53.8% 39.0% 6.8% 0.4%
Décor 58.4% 33.3% 6.5% 1.7%
Dishcloths 62.6% 27.5% 9.9%
Overall, the predominance of easy patterns reflects the accessibility of crochet as a craft and aligns with the goal of many project types to cater to a wide audience. The relative scarcity of experienced-level patterns suggests that most published resources emphasize broad usability rather than advanced expertise.
B.3 Pattern Complexity Analysis
B.3.1 Instruction Length Statistics
We first analyze the distribution of instruction lengths, measured in raw character counts. As shown in Table 13, the average instruction length is over 3,200 characters, while the median is substantially lower at 2,453 characters, reflecting a long-tailed distribution. The most complex patterns extend beyond 30,000 characters, while some very short patterns are as small as 20 characters.
Table 13: Instruction length statistics (in characters).
Statistic Value
Average 3216.0
Median 2453.0
Min 20
Max 30634
25th percentile 1511.8
75th percentile 4136.2
90th percentile 6403.9
Out of 6,085 total patterns, 6,084 (99.98%) contain full instructions.
B.3.2 Abbreviation Statistics
Abbreviations, such as sc, dc, and hdc, are a distinctive element of crochet instructions. Table 14 reports abbreviation counts across all patterns. Most patterns contain about 10 abbreviations, with values ranging from 1 to 31.
Table 14: Abbreviation count statistics.
Statistic Value
Average 10.6
Median 10.0
Min 1
Max 31
B.3.3 Complexity by Skill Level
Instruction length correlates with the designated skill level. As shown in Table 15, beginner-level patterns average under 2,000 characters, while intermediate patterns extend to over 4,200. Experienced patterns are the longest, averaging 7,689 characters. Rare categories such as easy to intermediate skew extremely long due to outliers.
Table 15: Instruction length and abbreviation counts by skill level.
Skill Level Avg. Length Median Length Avg. Abbr. Count
Easy to intermediate 13812.0 13812.0 21.0 1
Experienced 7689.4 6729.0 9.8 80
Intermediate 4221.3 3387.0 10.7 1967
Easy 2760.7 2182.0 10.8 3569
Beginner 1673.9 1365.0 9.2 465
Beginners 1633.0 1633.0 11.0 1
Beginner/Easy 1063.0 1063.0 1
B.3.4 Most and Least Complex Project Types
Finally, we identify the most complex and simplest project types by average instruction length. Tables 16 and 17 list the top 10 categories. Garments such as dresses, vests, pants, and tunics are the most demanding, with average instructions exceeding 5,800 characters. By contrast, smaller accessories such as cowls, washcloths, scarves, and headbands are substantially shorter, typically under 2,000 characters.
Table 16: Top 10 most complex project types (by average instruction length).
Project Type Avg. Length Median Count
Dresses 6484.9 5799.0 34
Vests 6032.0 5193.5 64
Pants 5866.7 5409.0 11
Tunics 5850.4 5832.0 29
Sets 5625.5 4847.0 111
Sweaters & Cardigans 5429.2 5113.0 357
Amigurumi & Toys 5322.4 4505.0 286
Jackets 5311.9 4831.0 31
Onesies & Rompers 5263.4 5181.0 5
Aprons 4467.8 4494.0 11
Table 17: Top 10 simplest project types (by average instruction length).
Project Type Avg. Length Median Count
Cowls 1288.3 956.5 154
Washcloths & Mitts 1502.5 1420.0 28
Scarves 1567.3 1221.0 304
Headbands 1617.5 1475.5 38
Dishcloths 1688.4 1571.0 222
Coasters 1750.3 1625.0 26
Booties 1921.9 1938.5 24
Jewelry 1960.3 1549.0 55
Super Scarves 2007.6 1213.0 13
Tech Accessories 2011.1 2099.0 13
Taken together, these results highlight strong alignment between project type, designated skill level, and instruction length. Garment-oriented projects require substantially longer and more complex instructions, while accessories and small decorative items remain simple and concise.
Appendix C Prompts
C.1 Task A: Stitch Recognition Prompt
This task evaluates a model’s ability to identify stitches present in a crochet product image.
Stitch Recognition Prompt (Rendered Example) SYSTEM PROMPT You are a crochet stitch expert. Given an image of a crochet product, identify all stitches that appear. Requirements: - Use only standard U.S. crochet abbreviations (e.g., sc, hdc, dc, tr, ch, sl st, pop, etc.). - Output must be a comma-separated list of abbreviations. - Do not include explanations, extra text, or formatting beyond the list. USER PROMPT Look at this crochet product image and list the stitches used. [Image]
C.2 Task B: Instruction Selection Prompt
This task evaluates a model’s ability to choose the correct instructions from multiple-choice options.
[…]
C.3 Task C: Instruction Generation Prompt
This task evaluates a model’s ability to generate complete crochet instructions from an image.
Instruction Generation Prompt (Rendered Example) SYSTEM PROMPT You are a professional crochet pattern writer. Examine the image of the finished crochet product carefully. Write a complete set of crochet instructions in the standard style used in published patterns. Requirements: - Use standard abbreviations: sc (single crochet), hdc (half double crochet), dc (double crochet), tr (treble), ch (chain), sl st (slip stitch), rep (repeat). - Organize the instructions row by row or round by round (e.g., ”Rnd 1: …”, ”Row 2: …”). - If color changes are visible in the image, include them in the pattern. - Keep the instructions concise and precise, as if for experienced crocheters. - Output only the crochet pattern. Do not add any explanations, commentary, or extra text. USER PROMPT Generate step-by-step crochet instructions for this image. [Image]
C.4 Task D (Step-level): NL $\rightarrow$ DSL Translation Prompt
This task evaluates whether a model can translate a single natural language instruction into exactly one line of compilable CrochetPARADE DSL code.
Step-level NL $\rightarrow$ DSL Translation Prompt (Rendered Example) SYSTEM PROMPT You are a crochet compiler. Translate the next instruction NL into one line of CrochetPARADE DSL. Use consistent naming and syntax. Important rules for translations: 1. Make sure your output ONLY contains the DSL code, nothing else. 2. Use the previous examples to understand the pattern of translation. 3. Be consistent in naming conventions with the examples. 4. Your output should be exactly one line of DSL code. USER PROMPT Now translate the NL into DSL: NL: DSL:
C.5 Task D (Project-Level): NL $\rightarrow$ DSL Translation Prompt
This task evaluates whether a model can convert natural language crochet instructions (with optional images) into compilable CrochetPARADE DSL code.
NL $\rightarrow$ DSL Translation Prompt (Rendered Example) SYSTEM PROMPT You are a professional crochet pattern writer. Convert instructions + images into compilable CrochetPARADE DSL code. Output only the DSL code. No explanations, commentary, or extra text. Example 1: ”image path”: https://www.yarnspirations.com/cdn/shop/files/BRC0116-035467M.jpg, INSTRUCTIONS Note: Join with sl st to first sc at end of each rnd. Ch 2. **Rnd 1:** 6 sc in 2nd ch from hook. Join. (6 sc) **Rnd 2:** Ch 1. 2 sc in each sc around. Join. (12 sc) **Rnd 3:** Ch 1. (2 sc in next sc, 1 sc in next sc) repeat around. End with 1 sc. Join. (18 sc) **Rnd 4:** Ch 1. (2 sc in next sc, 1 sc in each of next 2 sc) repeat. End with 1 sc in last 2 sc. Join. (24 sc) **Rnd 5:** Ch 1. Sc in each sc around. Join. (24 sc) **Rnd 6:** Ch 1. (2 sc in next sc, 1 sc in each of next 3 sc) repeat. End with 1 sc in last 3 sc. Join. (30 sc) **Rnds 7–8:** Repeat Rnd 5 (sc in each sc). Join. (30 sc each round) **Rnd 9:** Ch 1. **Working in back loops only**: (2 sc in next sc, 1 sc in each of next 2 sc) repeat. End with 1 sc in last 2 sc. Join. (40 sc) **Rnd 10:** Ch 1. Sc in each sc around (both loops). Join. (40 sc) **Rnd 11:** Ch 1. (2 sc in next sc, 1 sc in each of next 3 sc) repeat. End with 1 sc in last 3 sc. Join. (50 sc) **Finish:** Fasten off. DSL ¶ch.B ¶sc@B.A,5sc@B,ss@A ¶ch.A,sk,6sc2inc,ss@A ¶ch.A,sk,[sc2inc,sc]*6,ss@A ¶ch.A,sk,[sc2inc,2sc]*6,ss@A ¶ch.A,sk,24sc,ss@A ¶ch.A,sk,[sc2inc,3sc]*6,ss@A ¶[ch.A,sk,30sc,ss@A ¶]*2 ¶ch.A,sk,[scbl,scbl@[@],2scbl]*10,ss@A ¶ch.A,sk,40sc,ss@A ¶ch.A,sk,[sc2inc,3sc]*10,ss@A USER PROMPT Now generate DSL code for the following: [Image] [Instructions] Rnd 1: Ch 2, 6 sc in ring Rnd 2: 2 sc in each (12) Rnd 3: [Sc, sc, inc] around (16) Rnd 4: [Tr, sc] repeat around [DSL]
Appendix D CrochetPARADE: Pattern Renderer, Analyzer, and Debugger
CrochetPARADE (short for Crochet Pattern Renderer, Analyzer, and Debugger) is an interactive platform that enables users to author, visualize, test, and export crochet patterns in both 2D and 3D [crochetparade]. By combining a custom pattern grammar with simulation and rendering tools, CrochetPARADE addresses common issues of ambiguity, correctness, and interpretability in textual crochet instructions. [^4] https://www.crochetparade.org/
Core Capabilities.
  • Interactive authoring and rendering. Users write pattern instructions in the CrochetPARADE grammar and then invoke a “calculate” operation to convert those instructions into a virtual model. The system supports both 2D and 3D views, along with interactive controls such as zoom, rotation, and stitch highlighting.
  • Validation and debugging. CrochetPARADE parses the input, checks for syntactic and consistency errors (e.g., mismatched stitch counts, impossible attachments), and flags over- or under-stretched stitches.
  • Export and interoperability. From a rendered pattern, users can export:
    • A standard crochet chart (SVG) with conventional stitch symbols and labeled stitch connections.
    • A 3D model (GLTF format) for integration into external tools such as Blender.
    • The underlying pattern instructions text (in the CrochetPARADE grammar), ensuring reproducibility and sharing.
Design Ideals and Rationale.
CrochetPARADE is built to meet several design goals: (i) unambiguous precision, where the grammar is far more strict than free-form natural language, reducing interpretive errors; (ii) local computation, since all parsing, simulation, and rendering occur client-side in the browser with no user instructions sent to a central server; and (iii) open source extensibility, as the platform is released under GPLv3, with the grammar manual provided under a Creative Commons BY-NC-SA license.
Role in Our Work.
Within the context of CrochetBench, CrochetPARADE provides a rigorous target representation: model predictions can be compiled into CrochetPARADE instructions, validated for syntactic and structural correctness, and then visualized or executed. This enables evaluation beyond surface-level metrics (e.g., BLEU, ROUGE) toward executor correctness—whether a generated pattern is valid, renderable, and stitch-balanced.
Figure 3: Example of the CrochetBench translation pipeline. (Left) Natural language crochet instructions from the dataset. (Second) Automatically translated into CrochetPARADE DSL, a formal stitch grammar. (Third) Mesh rendering generated from the DSL. (Right) Target crocheted item image provided in the dataset. This pipeline enables direct text-to-image consistency checks, automated validation, and future training of NL $\rightarrow$ DSL models, analogous to text-to-code generation.
Appendix E DSL Error Taxonomy
To better understand failure cases in Task D, we extend the validator’s error analysis with detailed subcategories and examples. Errors fall into four major groups:
1. Syntax Structure Errors
Unbalanced Brackets.
Missing opening/closing parentheses or brackets.
Examples Unbalanced brackets: (sc,hc5,sltr)infl)
Multiple References Without Parentheses.
Improper formatting of references.
Example Multiple references defined without parenthesis: (21ch),turn
sk,(20sc)
(2ndrow):Ch1.(1scbl)ineachchtoendofrow.Turn
2. Stitch Definition Errors
Stitch Not Found.
Undefined stitch types not in the dictionary.
Examples ch1, ch3, scfp, hdc_bar
3. Labeling and Reference Errors
Label Not Found.
Reference to a non-existent label.
Example Label not found: C
Reusing Labels Incorrectly.
Same label used for non-adjacent stitches.
Example Cannot use same label over non-adjacent stitches. Consider using different labels.
4. Structural and Formatting Issues
Turning Errors.
Misplaced turning commands.
Example Turning can happen only at the end of a row.
Variable Naming Issues.
Conflicts between variable names and stitch names.
Example Error: variable name matches stitch name. For example, \$ch=0\$ cannot be used since ’ch’ is a stitch name.
Runtime Errors.
Low-level parsing failures from the JavaScript compiler.
Examples Cannot read properties of null (reading ’0’) Cannot use ’in’ operator to search for ’attach_id’ in NaN